### v1

@@ 5 Approximate-MMS with Copies: Ordered Instances
First, in Section 5.1, we review some of the known reduction rules and introduce new ones. Then, in Section 5.2, we present a simple algorithm $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ which combines the idea of bag-filling and round robin to achieve an $\alpha$-MMS allocation with copies (see Algorithm 4). This algorithm is the building block for achieving $6/7$-MMS and $4/5$-MMS with copies in the the sections that follow. In particular, in Sections 5.3 and 5.4, we obtain $6/7$-MMS with $\lfloor n/2\rfloor$ distinct copies and $4/5$-MMS with $\lfloor n/3\rfloor$ distinct copies respectively by first reducing the initial instance using appropriate valid reductions, and then applying $\mathsf{BagFill\mbox{-}RoundRobin}$.
For the rest of this section, we only consider ordered instances.
@@ 5.1 Reduction Rules
Rule $S^{\alpha}(\mathcal{I})$ Preconditions: $\mathcal{I}$ is ordered and $R^{\alpha}_{1}$ is not applicable. There exists an agent $i$ such that $v_{i}(\{1,n+1\})\geq\alpha\cdot\mu_{i}(\mathcal{I})$. Process: Allocate $\{1,n+1\}$ to $i$. Set $N^{\prime}\leftarrow N\setminus\{i\}$ and $M^{\prime}\leftarrow M\setminus\{1,n+1\}$. If there exists agent $j\neq i$ such that $v_{i}(\{1,n+1\})\geq\alpha\cdot\mu_{j}(\mathcal{I})$: - Duplicate good $1$ and allocate $\{1,n\}$ to agent $j$. - Set $N^{\prime}\leftarrow N^{\prime}\setminus\{j\}$ and $M^{\prime}\leftarrow M^{\prime}\setminus\{n\}$. Guarantees: $\mathcal{I}^{\prime}=(N^{\prime},M^{\prime},\{v_{i}\}_{i\in N^{\prime}})$ is ordered. For all $i\in N^{\prime}$, $\mu_{i}(\mathcal{I}^{\prime})\geq\mu_{i}(\mathcal{I})$.
@@
Rule $T^{\alpha}(\mathcal{I})$ Preconditions: $\mathcal{I}$ is ordered and $R^{\alpha}_{1}$ is not applicable. There exists three different agents $i,j,\ell$ such that $v_{i}(\{1,n+1\})\geq\alpha\cdot\mu_{i}(\mathcal{I})$, $v_{j}(\{2,n+2\})\geq\alpha\cdot\mu_{j}(\mathcal{I})$, and $v_{\ell}(\{1,n+3\})\geq\alpha\cdot\mu_{\ell}(\mathcal{I})$. Process: Allocate $\{1,n+1\}$ to $i$. Allocate $\{2,n+2\}$ to $j$. Duplicate good $1$ and allocate $\{1,n+3\}$ to $\ell$. Set $N^{\prime}\leftarrow N\setminus\{i,j,\ell\}$ and $M^{\prime}\leftarrow M\setminus\{1,2,n+1,n+2,n+3\}$. Guarantees: $\mathcal{I}^{\prime}=(N^{\prime},M^{\prime},\{v_{i}\}_{i\in N^{\prime}})$ is ordered. For all $i\in N^{\prime}$, $\mu_{i}(\mathcal{I}^{\prime})\geq\mu_{i}(\mathcal{I})$.
@@
For all agents $i$ who is assigned a bundle $B$ in $T^{\alpha}$, we have $v_{i}(B)\geq\alpha\cdot\mu_{i}(\mathcal{I})$. For all the remaining agents $i$, we have $\alpha\cdot\mu^{n}_{i}(M)\geq v_{i}(\{1\})\geq v_{i}(\{2\})$ and $\alpha/2\cdot\mu^{n}_{i}(M)\geq v_{i}(\{2n+1\})\geq v_{i}(2n+2)\geq v_{i}(2n+3)$. Since $\alpha\leq 4/5$, by Lemma 11, $v^{n-3}_{i}(M\setminus\{1,2,n+1,n+2,n+3\})\geq v^{n}_{i}(M)$. ∎
Definition 5.
We call an ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, $R^{\alpha}$-irreducible, if for all $k<|M|/|N|$, $R^{\alpha}_{k}(\mathcal{I})$ is not applicable. Similarly, an ordered instance is $S^{\alpha}$-irreducible and $T^{\alpha}$-irreducible, if $S^{\alpha}$ and $T^{\alpha}$ are not applicable respectively.

### v2

@@ 5 Approximate-MMS with Copies: Additive Valuations
In Section 5.2, we review some of the known reduction rules and introduce new ones. Then, in Section 5.3, we present a simple algorithm $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ which combines the idea of bag-filling and round robin to achieve an $\alpha$-MMS allocation with copies (see Algorithm 4). This algorithm is the building block for achieving $6/7$-MMS and $4/5$-MMS with copies in the sections that follow. In particular, in Sections 5.4 and 5.5, we obtain $6/7$-MMS with $\lfloor n/2\rfloor$ distinct copies and $4/5$-MMS with $\lfloor n/3\rfloor$ distinct copies, respectively, by first reducing the initial instance using appropriate valid reductions, and then applying $\mathsf{BagFill\mbox{-}RoundRobin}$.
5.1 Ordered Instance
Given an instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, for every agent $i\in N$, let $\sigma_{i}:[m]\to[m]$ be a permutation that orders the goods according to $i$’s preferences, namely $\sigma_{i}(j)$ is the $j$th most valued good (breaking ties by index111any consistent tie breaking rule will do., and thus $v_{i}(\sigma_{i}(1))\geq v_{i}(\sigma_{i}(2))\geq\dots\geq v_{i}(\sigma_{i}(m))$.
As done by Bouveret and Lemaître [16], we construct an ordered instance $\mathcal{I}^{\prime}=(N,M^{\prime},\{v^{\prime}_{i}\}_{i\in N})$ with a new set of “virtual goods”, $M^{\prime}=\{1,\dots,m\}$, and modified valuations, $\{v^{\prime}_{i}\}_{i\in N}$, such that $v^{\prime}_{i}(j)=v_{i}(\sigma_{i}(j))$ for any $j\in M^{\prime}$.
Definition 4 (Ordered Counterpart ([16])).
Given a fair division instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, its ordered counterpart, $\mathcal{I}^{\prime}=(N,M^{\prime},\{v^{\prime}_{i}\}_{i\in N})$, is an instance with the same set of agents and a set of “virtual goods”, $M^{\prime}$, of the same cardinality as $M$. For each agent $i\in[n]$, and any $j\in M^{\prime}$, it holds that $v^{\prime}_{i}(j)=v_{i}(\sigma_{i}(j))$.
Observe that by definition, all agent in $\mathcal{I}^{\prime}$ have the same ordinal preferences, i.e., for every $i\in N$, $v^{\prime}_{i}(1)\geq v^{\prime}_{i}(2)\geq\dots\geq v^{\prime}_{i}(m)$. It was shown in [16] that to guarantee agent $i$ a value of at least $\alpha_{i}$ in an instance $\mathcal{I}$, it suffices to find an allocation providing this guarantee in its ordered counterpart $\mathcal{I}^{\prime}$. This is cast in the following proposition.
Proposition 2 (Reduction to ordered without copies [16]).
Let $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ be a fair division instance and let $\mathcal{I}^{\prime}=(N,M^{\prime},\{v^{\prime}_{i}\}_{i\in N})$ be its ordered counterpart. Any allocation (without copies) $A^{\prime}=(A^{\prime}_{1},\dots,A^{\prime}_{n})$ for $\mathcal{I}^{\prime}$ can be converted in poly-time to an allocation (without copies) $A=(A_{1},\dots,A_{n})$ for $\mathcal{I}$, such that for every agent $i$, $v_{i}(A_{i})\geq v^{\prime}_{i}(A^{\prime}_{i})$.
Proof.
The transformation from the ordered allocation $A^{\prime}$ to the allocation $A$ is done via a simple picking sequence procedure: In every iteration $j\in\{1,\dots,m\}=M^{\prime}$, let $i$ be the agent such that $j\in A^{\prime}_{i}$. Agent $i$ chooses the most-valuable item in $M$ (according to $v_{i}$), which has not yet been chosen.
Since by iteration $j$ exactly $j-1$ items have been chosen, agent $i$ always chooses a good $g\in M$ such that $v_{i}(g)\geq v_{i}(\sigma_{i}(j))=v^{\prime}_{i}(j)$ and thus, by additivity, $v_{i}(A_{i})\geq v^{\prime}_{i}(A^{\prime}_{i})$. ∎
We will show a similar result to Proposition 2, for allocations with $t$ distinct copies in which the top $t$ goods (and their copies) are allocated to $2t$ different agents.
Definition 5 (Simple Allocation).
Given an ordered instance $\mathcal{I}^{\prime}=(N,M^{\prime},\{v^{\prime}_{i}\}_{i\in N})$, an allocation with $t$ total copies $A^{\prime}=(A^{\prime}_{1},\dots,A^{\prime}_{n})$ is called simple if
  • There is at most one copy of each good.
  • The duplicated items are $g_{1},\dots,g_{t}$ for $t\leq n/2$.
  • There exists a set of $2t$ distinct agents, $i_{1},\dots,i_{t},i^{\dagger}_{1},\dots,i^{\dagger}_{t}$ such that for any $j\in[t]$, agents $i_{j}$ and $i^{\dagger}_{j}$ are allocated the two instances of $g_{j}$, i.e., $A^{\prime}_{i_{j}}\cap A^{\prime}_{i^{\dagger}_{j}}=\{g_{j}\}$.
Recall that agent $i$ cannot get two instances of the same good in $A^{\prime}_{i}$. [^2] Otherwise, and if additive valuations are extended so that two instances of the same good yield twice its value, then the same argument as in Proposition 2 would have implied the correctness of Proposition 3.
Proposition 3 (Reduction to ordered for simple allocations).
Let $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ be a fair division instance and let $\mathcal{I}^{\prime}=(N,M^{\prime},\{v^{\prime}_{i}\}_{i\in N})$ be its ordered counterpart. Let $A^{\prime}=(A^{\prime}_{1},\dots,A^{\prime}_{n})$ be a simple allocation for $\mathcal{I}^{\prime}$ with $t$ distinct copies. Then, $A^{\prime}$ can be converted in poly-time to an allocation for $\mathcal{I}$ with $t$ distinct copies, $A=(A_{1},\dots,A_{n})$, such that for every agent $i$, $v_{i}(A_{i})\geq v^{\prime}_{i}(A^{\prime}_{i})$.
In order to prove Proposition 3, we presented the following useful lemma.
Lemma 7.
Fix an agent $i\in N$ and a threshold $\tau\in\mathbb{R}_{+}$. Let $X,Y\subset M$ be such that
  • $|X|=|Y|=k$, $X=\{x_{1},\ldots,x_{k}\}$, $Y=\{y_{1},\ldots,y_{k}\}$, and
  • for all $j\in[k]$, $v_{i}(y_{j})\geq v_{i}(x_{j})$.
If there exists a set $G\subset M\setminus X$ of $\ell$ distinct goods such that $v_{i}(X\cup\{g\})\geq\tau$ for all $g\in G$, then there exists a set $H\subset M\setminus Y$ of $\ell$ distinct goods such that $v_{i}(Y\cup\{h\})\geq\tau$ for all $h\in H$.
Proof.
Observe that by additivity, any good $g\in G\setminus Y$ satisfies $v_{i}(Y\cup\{g\})\geq v_{i}(X\cup\{g\})\geq\tau$. Thus, it suffices to show that for every $g\in G\cap Y$, there exists a (distinct) corresponding good $h(g)\in M\setminus(Y\cup G)$ such that $v_{i}(Y\cup\{h(g)\})\geq\tau$. If this requirement is met, the set $H=(G\setminus Y)\cup\{h(g)\mid g\in G\cap Y\}$ satisfies the desired condition in the lemma. We construct the mapping from every $g$ to $h(g)$ using an augmenting path argument in a bipartite graph (see Figure 1 for an illustration).
Consider the directed bipartite graph $\mathcal{G}=(\mathcal{X},\mathcal{Y},\mathcal{E})$, where $\mathcal{X}=\{\mathcal{X}_{g}\}_{g\in X}$ and $\mathcal{Y}=\{\mathcal{Y}_{g}\}_{g\in Y}$. Every vertex $\mathcal{X}_{g}\in\mathcal{X}$ (respectively, $\mathcal{Y}_{g}\in\mathcal{Y}$ ) corresponds to good $g\in X$ (resp., $g\in Y$). Note that $|\mathcal{X}|=|\mathcal{Y}|=k$. When clear in the context, we abuse notation and write $\mathcal{X}_{j}$ instead of $\mathcal{X}_{x_{j}}$ (and similarly for $\mathcal{Y}$).
For every $j\in[k]$, we add the (black) edge $(\mathcal{Y}_{j},\mathcal{X}_{j})$. In addition, for any good $g\in X\cap Y$, we add the (red) edge $(\mathcal{X}_{g},\mathcal{Y}_{g})$. Observe that $\mathcal{E}$ is a collection of disjoint cycles of length 2, or paths333 Assume that both $X$ and $Y$ are ordered according to $i$’s valuation: $v_{i}(x_{1})\geq\dots\geq v_{i}(x_{k})$ and the same for $Y$, where tie breaking is consistent among $X$ and $Y$. Observe that path between $\mathcal{Y}_{y_{j}}$ and $\mathcal{Y}_{y_{j^{\prime}}}$ implies that agent $i$ prefers $y_{j}$ over $y_{j^{\prime}}$, so a cycle of length greater than 2 is not possible. . Additionally, every vertex $v\in\mathcal{X}\cup\mathcal{Y}$ satisfies $deg(v)\in\{1,2\}$, where by $deg$ we consider both in- and out-degrees.
Fix $g\in G\cap Y$. Since $g\in G$, then $g\notin X$ and $deg(\mathcal{Y}_{g})=1$. Let $P=(\mathcal{Y}_{g},\mathcal{X}_{g_{1}},\mathcal{Y}_{g_{1}},\dots,\mathcal{X}_{g_{\ell}},\mathcal{Y}_{g_{\ell}},\mathcal{X}_{g^{\prime}})$ be the unique path originating at $\mathcal{Y}_{g}$. We claim that $h(g)=g^{\prime}$ satisfies the desired requirement.
Since $\mathcal{X}_{g^{\prime}}$ has no outgoing edges, $g^{\prime}\notin Y$. Also, since $g^{\prime}\in X$, then $g^{\prime}\notin G$. Thus, $g^{\prime}\in M\setminus(Y\cup G)$. It remains to show that $v_{i}(Y\cup\{g^{\prime}\})\geq\tau$. To this end, let $Q=\{g_{1},\dots,g_{\ell}\}\subseteq X\cap Y$, as defined by the path $P$. Observe that the set of goods which correspond to $\mathcal{Y}\cap P$ is exactly $Q\cup\{g\}$ (see Figure 1). Similarly, the set of goods corresponding to $\mathcal{X}\cap P$ is exactly $Q\cup\{g^{\prime}\}$. Since every vertex in $\mathcal{Y}\setminus P$ is pointed to by exactly one vertex in $\mathcal{X}\setminus P$, and $v_{i}(y_{j})\geq v_{i}(x_{j})$ for every $j\in[k]$, it follows that $v_{i}(Y\setminus(\{Q\cup\{g\}))\geq v_{i}(X\setminus(\{Q\cup\{g^{\prime}\}))$. By the last inequality and additivity of $v_{i}$ we get
$v_{i}(Y\cup\{g^{\prime}\})$ $=v_{i}(Y\setminus(Q\cup\{g\}))+v_{i}(Q\cup\{g\})+v_{i}(\{g^{\prime}\})$
$\geq v_{i}(X\setminus(Q\cup\{g^{\prime}\}))+v_{i}(Q\cup\{g^{\prime}\})+v_{i}(\{g\})$
$=v_{i}(X\cup\{g\})$
$\geq\tau,$
which concludes the proof.
[…]
Figure 1: Illustration of the augmenting path argument in the proof of Lemma 7. The element $g\in G$ (whose corresponding node is $\mathcal{Y}_{g}$, $g=y_{1}$) is replaced with $g^{\prime}$ (whose corresponding node is $\mathcal{X}_{g^{\prime}}=\mathcal{X}_{4}$, $g^{\prime}=x_{4}$). We write $\mathcal{X}_{i}$ / $\mathcal{Y}_{i}$ rather than the cumbersome $\mathcal{X}_{x_{i}}$/$\mathcal{Y}_{y_{i}}$. It is purely an example that the path nodes are associated with a contiguous prefix of the goods, e.g., $\mathcal{X}_{1}$ could point to $\mathcal{Y}_{5}$.
Proof of Proposition 3.
Let $A^{\prime}=(A^{\prime}_{1},\dots,A^{\prime}_{n})$ be a simple allocation in which the goods $g_{1}<\ldots<g_{t}$ for $t\leq n/2$ are copied. Rename the agents so that for every $j\in[t]$ one copy of $g_{j}$ is allocated to agent $j$, i.e., $g_{j}\in A_{j}$. We will refer to each such $g_{j}$ as the ‘copy’, and to the other instance of $g_{j}$ (assigned to some agent $i\in N\setminus\{1,\dots,t\}$), as the ‘original’.
We will use a similar picking sequence procedure to the one given in Proposition 2, but with two rounds. First, we will only use the $m$ ‘original’ goods to determine the picking sequence, this is identical to the picking sequence presented earlier. Then, another picking sequence will use the ‘copies’ (and so will involve agents $1,\dots,t$, where each picks one good). Each agent in the second round may pick any good in $M$ which was not yet allocated in this round and was not picked by him in the first round. Let $A_{i}\subseteq M$ be the set of goods picked by agent $i$ in both rounds.
Any agent $i\in N\setminus\{1,\dots,t\}$ only participates in the first round, so as in Proposition 2, it holds that $v_{i}(A_{i})\geq v^{\prime}_{i}(A^{\prime}_{i})$. For agent $j\in\{1,\dots,t\}$, let $B_{j}$ be the set of $|A_{j}|-1$ goods he picked in the first round. To show that $v_{j}(A_{j})\geq v^{\prime}_{j}(A^{\prime}_{j})$, it is enough to show that at the beginning of the second round there is a set of goods $H\subseteq M\setminus B_{j}$, such that (i) $|H|\geq j$, (ii) each $h\in H$ satisfies $v_{j}(B_{j}\cup\{h\})\geq v^{\prime}_{j}(A^{\prime}_{j})$. This would imply that in round $j$, player $j$ can pick a good $h$ such that $A_{j}=B_{j}\cup\{h\}$ satisfies the claim.
Denote $A^{\prime}_{j}=\{j,j_{1},\dots,j_{k}\}$, $B_{j}=\{b_{1},\dots,b_{k}\}$, where $b_{\ell}$ was picked by agent $j$ in round $j_{\ell}$. The existence of $H$ follows from applying Lemma 7 with $X=\{\sigma_{j}(j_{1}),\dots,\sigma_{j}(j_{k})\}$, $Y=B_{j}$, $G=\{\sigma(g_{1}),\dots,\sigma_{j}(g_{j})\}$ and $\tau=v^{\prime}_{j}(A^{\prime}_{j})$. Observe that by definition of the first picking sequence, for any $\ell$, it holds that $v_{i}(b_{\ell})\geq v_{i}(\sigma_{j}(j_{\ell}))=v^{\prime}_{j}(j_{\ell})$. Also, $\tau=v^{\prime}_{j}(A^{\prime}_{j})=v^{\prime}_{j}(A^{\prime}_{j}\setminus\{g_{j}\})+v^{\prime}_{j}(g_{j})=v_{j}(X)+v_{j}(\sigma_{j}(g_{j}))=v_{j}(X\cup\{\sigma_{j}(g_{j})\})$. By definition, any good in $g\in G$ has $v_{j}(\{g\})\geq v_{j}(\sigma_{j}(g_{j}))$, and by additivity all the conditions of Lemma 7 are satisfied. This concludes the proof. ∎
For the rest of this section, we only consider ordered instances. Because the algorithms we use to solve these instances always output a simple allocation, and therefore, their guarantees, by Proposition 3, can be extended to any instance, even unordered.
@@ 5.2 Reduction Rules
Definition 7.
We call an ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, $R^{\alpha}_{k}$-irreducible, if $R^{\alpha}_{k}(\mathcal{I})$ is not applicable. Moreover, we call an ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, $R^{\alpha}$-irreducible, if for all $k<|M|/|N|$, $R^{\alpha}_{k}(\mathcal{I})$ is not applicable.
Observation 1.
If $\,\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ is ordered and $R^{\alpha}_{k}$-irreducible for some $k<|M|/n$, then $v_{i}(nk+1)<\alpha\cdot\mu_{i}/(k+1)$ for all agents $i\in N$.
Proof.
By irreducibility of $R^{\alpha}_{k}$ for $K=\{k(n-1)+1,k(n-1)+2,\ldots,nk,nk+1\}$, we have $v_{i}(K)<\alpha\cdot\mu_{i}$ for all $i\in N$. Since $v_{i}(nk+1)\leq v_{i}(nk)\leq\ldots\leq v_{i}(k(n-1)+1)$, we get $v_{i}(nk+1)\leq v_{i}(K)/(k+1)\leq\alpha\cdot\mu_{i}/(k+1)$. ∎
@@
Rule $S^{\alpha}(\mathcal{I})$ Preconditions: $\mathcal{I}$ is ordered and $R^{\alpha}_{1}$-irreducible. There exists an agent $i$ such that $v_{i}(\{1,n+1\})\geq\alpha\cdot\mu_{i}(\mathcal{I})$. Process: Allocate $\{1,n+1\}$ to $i$. Set $N^{\prime}\leftarrow N\setminus\{i\}$ and $M^{\prime}\leftarrow M\setminus\{1,n+1\}$. If there exists agent $j\neq i$ such that $v_{i}(\{1,n+1\})\geq\alpha\cdot\mu_{j}(\mathcal{I})$: - Duplicate good $1$ and allocate $\{1,n\}$ to agent $j$. - Set $N^{\prime}\leftarrow N^{\prime}\setminus\{j\}$ and $M^{\prime}\leftarrow M^{\prime}\setminus\{n\}$. Guarantees: $\mathcal{I}^{\prime}=(N^{\prime},M^{\prime},\{v_{i}\}_{i\in N^{\prime}})$ is ordered. For all $i\in N^{\prime}$, $\mu_{i}(\mathcal{I}^{\prime})\geq\mu_{i}(\mathcal{I})$.
@@
Rule $T^{\alpha}(\mathcal{I})$ Preconditions: $\mathcal{I}$ is ordered and $R^{\alpha}_{k}$-irreducible for $k\leq 1$. There exists three different agents $i,j,\ell$ such that $v_{i}(\{1,n+1\})\geq\alpha\cdot\mu_{i}(\mathcal{I})$, $v_{j}(\{2,n+2\})\geq\alpha\cdot\mu_{j}(\mathcal{I})$, and $v_{\ell}(\{1,n+3\})\geq\alpha\cdot\mu_{\ell}(\mathcal{I})$. Process: Allocate $\{1,n+1\}$ to $i$. Allocate $\{2,n+2\}$ to $j$. Duplicate good $1$ and allocate $\{1,n+3\}$ to $\ell$. Set $N^{\prime}\leftarrow N\setminus\{i,j,\ell\}$ and $M^{\prime}\leftarrow M\setminus\{1,2,n+1,n+2,n+3\}$. Guarantees: $\mathcal{I}^{\prime}=(N^{\prime},M^{\prime},\{v_{i}\}_{i\in N^{\prime}})$ is ordered. If $\alpha\leq 4/5$, for all $i\in N^{\prime}$, $\mu_{i}(\mathcal{I}^{\prime})\geq\mu_{i}(\mathcal{I})$.
@@
For all agents $i$ who is assigned a bundle $B$ in $T^{\alpha}$, we have $v_{i}(B)\geq\alpha\cdot\mu_{i}(\mathcal{I})$. Note that by the preconditions of the reduction $T^{\alpha}(\mathcal{I})$, $\mathcal{I}$ is $R^{\alpha}_{k}$-irreducible for $k\leq 1$. Thus, by Observation 1, for all the remaining agents $i$, we have $\alpha\cdot\mu^{n}_{i}(M)\geq v_{i}(1)\geq v_{i}(2)$ and $\alpha/2\cdot\mu^{n}_{i}(M)\geq v_{i}(n+1)\geq v_{i}(n+2)\geq v_{i}(n+3)$. Since $\alpha\leq 4/5$, by Lemma 12, $v^{n-3}_{i}(M\setminus\{1,2,n+1,n+2,n+3\})\geq v^{n}_{i}(M)$. ∎
Remark 1.
In subsequent sections, similar to Definition 7, an ordered instance is $S^{\alpha}$-irreducible and $T^{\alpha}$-irreducible, if $S^{\alpha}$ and $T^{\alpha}$ are not applicable respectively.